# Reasoning-Enhanced Object-Centric Learning for Videos

### v1

@@ Appendix F STATM Structure in Different Model
Due to the differences between the SAVi/SAVi++ and STEVE models, there are certain distinctions in the enhanced models’ STATM module as well. The following is a simplified algorithm for the STATM module in different models:
Table 10: Simplified algorithm for the STATM module in different models.
@@
  Input: $S_{t}$, $M_{t}$
  $X_{t}$ = Spatiotemporal Attention($S_{t}$, $M_{t}$, $M_{t}$)
  $X_{t}$ = LayerNorm($X_{t}$ + $S_{t}$)
  $Y_{t}$ = MLP($X_{t}$)
  $Y_{t}$ = LayerNorm($Y_{t}$)
  Return: $X_{t}$ + $Y_{t}$
  Input: $S_{t}$, $M_{t}$
  $S_{t}$ = LayerNorm($S_{t}$)
  $M_{t}$ = LayerNorm($M_{t}$)
  $X_{t}$ = Spatiotemporal Attention($S_{t}$, $M_{t}$, $M_{t}$)
  $X_{t}$ = LayerNorm($X_{t}$ + $S_{t}$)
  $Y_{t}$ = MLP($X_{t}$)
  Return: $S_{t}$ + $Y_{t}$

### v2

@@ Appendix F STATM Structure in Different Model
Due to the differences between the SAVi/SAVi++ and STEVE models, there are certain distinctions in the enhanced models’ STATM module as well. Table S8 illustrates the simplified algorithm for the STATM module in different models.
Table S8. Simplified algorithm for the STATM module in different models.
@@
  Input: $S_{t}$, $M_{t}$         $X_{t}$ = Spatiotemporal Attention($S_{t}$, $M_{t}$, $M_{t}$)   $X_{t}$ = LayerNorm($X_{t}$ + $S_{t}$)   $Y_{t}$ = MLP($X_{t}$)   $Y_{t}$ = LayerNorm($Y_{t}$)   Return: $X_{t}$ + $Y_{t}$   Input: $S_{t}$, $M_{t}$   $S_{t}$ = LayerNorm($S_{t}$)   $M_{t}$ = LayerNorm($M_{t}$)   $X_{t}$ = Spatiotemporal Attention($S_{t}$, $M_{t}$, $M_{t}$)   $X_{t}$ = LayerNorm($X_{t}$ + $S_{t}$)   $Y_{t}$ = MLP($X_{t}$)      Return: $S_{t}$ + $Y_{t}$
@@
Video
G.T.
SAVi++
Ours++
Video
G.T.
SAVi++
Ours++
Figure S1. Qualitative results of our model compared to SAVi++ on the MOVi-A dataset.
Video
G.T.
SAVi++
Ours++
Video
G.T.
SAVi++
Ours++
Figure S2. Qualitative results of our model compared to SAVi++ on the MOVi-B dataset.
Video
G.T.
SAVi++
Ours++
Video
G.T.
SAVi++
Ours++
Figure S3. Qualitative results of our model compared to SAVi++ on the MOVi-C dataset.
Video
G.T.
SAVi++
Ours++
Video
G.T.
SAVi++
Ours++
Figure S4. Qualitative results of our model compared to SAVi++ on the MOVi-D dataset.
Video
G.T.
SAVi++
Ours++
Video
G.T.
SAVi++
Ours++
Figure S5. Qualitative results of our model compared to SAVi++ on the MOVi-E dataset.
Figure S6. Attention map visualization on the MOVi-A dataset.
Figure S7. Attention map visualization on the MOVi-B dataset.
Figure S8. Attention map visualization on the MOVi-C dataset.
Figure S9. Attention map visualization on the MOVi-D dataset.
Figure S10. Attention map visualization on the MOVi-E dataset.